If a is equal to 1.2 ± 0.1cm, b=2.0 ± 0.2cm, c=1.05±0.05cm, calculate the following with the errors and standard errors


(i) ab/C³


(ii) ABC


(iii) C²a³/b²

1
Expert's answer
2022-02-20T15:43:35-0500

1)

Angular frequency "w=\frac{1}{\sqrt{LC}}"

Put value

"w=\frac{1}{\sqrt{8\times10^{-3}\times12.5\times10^{-6}}}=3162.27rad/sec"

Impedance at resonance frequency

"Z_{max}=R=175\Omega"

Maximum current

"I=\frac{V}{R}=\frac{25}{175}=0.142A"

"I_L=\frac{V}{X_L}=\frac{25}{25.29}=0.98A"

Potential difference

"V_c=I\times X_L"

"V_c=0.142\times25.29=3.59V"



"V_R=IR=0.142\times175=24.85V"

"V_c=0.142\times25.29=3.59V"